OpenStudy (anonymous):

How do I complete the square of: 9x²+36x-4y²-16y=16

OpenStudy (anonymous):

If I factor: 9(x²+4x)-4(y²-4y)=16 then what

OpenStudy (anonymous):

complete the squares in the parenthesis

OpenStudy (anonymous):

How do I manipulate the outside numbers afterward?

OpenStudy (anonymous):

9(x²+4x+4)-4(y²-4y+4)=16

OpenStudy (anonymous):

gimme a sec to type plz

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

\[9(x^{2}+4x+4)-4(y^{2}-4y+4)=16\]

OpenStudy (anonymous):

\[9(x^{2}+4x+4)-4(y^{2}-4y+4)=16+36-16\]

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

oh ok yeah, you have to multiply by the outside number

OpenStudy (anonymous):

thanks a lot youve been a big help

OpenStudy (anonymous):

Np.